A 75.0 g rock is dropped into a graduated cylinder containing 75.0 mL of water.  After the rock is dropped in, the water level rises to 78.1 mL.  What is the density of the rock?


1
Expert's answer
2021-09-29T04:59:47-0400

78.1-75= 3.1

D = M/V

= 75/3.1 = 24.194g/ml